# Hybrid-Emba3D: Geometry-Aware and Cross-Path Feature Hybrid Enhanced State Space Model for Point Cloud Classification

### v1

@@ 5 Conclusion
In this paper, we present Hybrid-Emba3D, a novel mamba-based Architecture for point cloud classification task, designing two key components: the geometry-feature coupling mechanism that aggregates geometric information from neighboring points around each centroid and central feature to enhance local geometric representation, and A multidimensional feature hybrid enhancer employing dual-path Architecture to strengthen global context modeling for BiSSM. Notably, the total computational cost with the two components is very small. Especially in the operations of coupling local geometric information into central features in a parameter-free manner, incurring no additional parameters, that is completely driven by local neighborhoods’ intrinsic geometric relationships.
Extensive experiments demonstrate Hybrid-Emba3D’s superior performance across classification benchmarks, particularly achieving new SOTA accuracy on ModelNet40 at the cost of minimal parameters. These also show that the Mamba architecture has potential, yet insufficient exploration in point cloud processing.

### v2

@@ 5 Conclusion
In this paper, we present HyMamba, a novel mamba-based Architecture for the point cloud classification task, designing two key components: the geometry-feature coupling mechanism that aggregates geometric information from neighboring points around each centroid and central feature to enhance local geometric representation, and a multidimensional feature hybrid enhancer employing the dual-path architecture to strengthen global context modeling for BiSSM. Notably, the total computational cost with the two components is very small. Especially in the operations of coupling local geometric information into central features in a parameter-free manner, which is completely driven by local neighborhoods’ intrinsic geometric relationships.
Extensive experiments demonstrate HyMamba’s superior performance across classification benchmarks, particularly achieving new SOTA accuracy on ModelNet40 at the cost of minimal parameters. These also show that the Mamba architecture has potential, yet insufficient exploration in point cloud processing. Future research could dynamically adjust the point cloud serialization strategy or SSM scanning strategy based on the relationships between points, so that the scanning path can reflect the spatial relationships between points to compensate for its lack of directness in modeling 3D spatial structures. Additionally, integrating CNN architectures or optimizing SSM’s long-sequence processing could enhance the representation of hierarchical local features.
Acknowledgements
All authors thank the 512 Lab and 513 Lab of the School of Weapon Science and Technology at Xi’an Technological University.
CRediT authorship contribution statement
Bin Liu: Conceptualization, Data curation, Formal analysis, Methodology, Project administration, Resources, Software, Writing – original draft. Chunyang Wang: Funding acquisition, Supervision. Xuelian Liu: Visualization, Validation. Bo Xiao: Investigation, Supervision. Guan Xi: Supervision,.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.